# Cut-based Conflict Analysis in Mixed Integer Programming

### v1

@@ 4 Reduction Techniques for Pure Binary Programs
Algorithm 1 shows the base algorithm for all variants of cut-based conflict analysis considered in this paper. The algorithm is initialized with an infeasible state $\rho_{p,q}$ and a conflicting constraint $C_{\textnormal{confl}}$ in $\rho_{p,q}$. First, the learned conflict constraint $C_{\textnormal{learn}}$ is set to the conflict constraint $C_{\textnormal{confl}}$. In each iteration, we extract the variable $x_{r}$ whose bound was changed in $\rho_{p,q}$. If the bound change was due to propagation of a constraint, then we extract the reason constraint $C_{\textnormal{reason}}$ that propagated $x_{r}$. In line 1 we “reduce” the reason constraint such that the resolvent, i.e., linear combination, of $C_{\textnormal{learn}}$ and the reduced reason $C_{\textnormal{reason}}$ (Line 1) that cancels $x_{r}$ remains infeasible. The learned conflict constraint is set to the resolvent, which can then be strengthened in Line 1 by, e.g., applying coefficient tightening as in SAT resolution conflict analysis. Finally, the state $\rho_{p,q}$ is set to the smallest state, with respect to the lexicographic order, such that the resolvent is infeasible. We continue this process until we reach an FUIP ($C_{\textnormal{learn}}$ is asserting) or $C_{\textnormal{learn}}$ proves global infeasibility.
@@ 9             $C_{\textnormal{learn}}\leftarrow\textnormal{strengthen}(C_{\textnormal{learn}})$
10
12return $C_{\textnormal{learn}}$
@@
The conflict analysis algorithm can fail if the resolvent becomes feasible. As explained in the previous section, this occurs if and only of the reason constraint $C_{\textnormal{reason}}$ does not propagate tightly. In the next sections, we present various reduction algorithms for the reason constraint $C_{\textnormal{reason}}$ that ensure tight propagation: the SAT-like clausal-based reduction Achterberg (2007a), which reduces the reason constraint to a clause responsible for the propagation; the coefficient tightening-based reduction, which is a direct generalization of Chai and Kuehlmann (2005b); and a stronger version of the MIR-based reduction presented in Mexi et al. (2023).
@@
The reason constraint is of the form
$C_{\textnormal{reason}}:x_{r}+\sum_{j\in J}a_{j}x_{j}\geq b,$
where $r\not\in J\subseteq\mathcal{I}$ and $a_{j}\geq 0$ for all $j\in J$. Furthermore, $C_{\textnormal{reason}}$ propagates $x_{r}$ in $\rho$ non-tightly, i.e.,
$b-\max\sum_{j\in J}a_{j}x_{j}\,\not\in\,\mathbb{Z}$
@@
4.1 Clause-based Reduction
Let $C_{\textnormal{reason}}$ be as in Assumption 1. In the graph-based conflict analysis as developed for MIP by Achterberg (2007a), the reason constraint is reduced to a clause that propagates $x_{r}$ tightly. This clause contains $x_{r}$ and all variables from the reason that are responsible for the propagation of $x_{r}$, namely $x_{r}\bigvee_{j\in J:u_{j}^{\rho}=0}x_{j}$, which corresponds to the linear constraint
$C_{\textnormal{clausal}}\,:x_{r}+\sum_{j\in J:u_{j}^{\rho}=0}x_{j}\geq 1.$
Note that this corresponds to the well-known cover cuts for knapsack constraints.
@@ 4.2 Coefficient Tightening-based Reduction
algorithm 2 summarizes the coefficient tightening-based reduction. Similar to the implementation in a PB solver Chai and Kuehlmann (2005a), in each iteration, the algorithm picks a relaxable variable in the reason constraint different from the variable we are resolving on and weakens it. Then it applies coefficient tightening to the resulting constraint. After each iteration, the algorithm checks if the resolvent of the reason constraint and conflict constraint is infeasible. In this case, the algorithm terminates and returns the reduced reason constraint.
Input : conflict constraint $C_{\textnormal{confl}}$, reason constraint $C_{\textnormal{reason}}$,
variable to resolve $x_{r}$, state $\rho$
@@
2       $x_{s}\leftarrow$ relaxable variable in $C_{\textnormal{reason}}\backslash\{x_{r}\}$
@@ 5      $C_{\textnormal{reason}}\leftarrow\textnormal{coefTight}(C_{\textnormal{reason}})$
Next, we prove that the reduction algorithm ensures that the key requirement of tight propagation holds at the latest after weakening all relaxable variables.
Proposition 2.
Let $C_{\textnormal{reason}}$ be as in Assumption 1. After weakening all variables in $P=\{j\in J\,:\,u_{j}^{\rho}=1\}$ and applying coefficient tightening, the resulting constraint propagates $x_{r}$ tightly.
Proof.
First, we can rewrite the constraint $C_{\textnormal{reason}}$ as
$x_{r}+\sum_{j\notin P}a_{j}x_{j}+\sum_{j\in P}a_{j}x_{j}\geq b.$
After weakening all variables from $P$, the constraint $C_{\textnormal{reason}}$ becomes
$x_{r}+\sum\limits_{j\notin P}a_{j}x_{j}\geq b-\sum\limits_{j\in P}a_{j}=:\tilde{b}.$ (10)
By assumption all variables not in $P$ are fixed to 0 and $C_{\textnormal{reason}}$ propagates $x_{r}$ non-tightly. Therefore, it must hold that $10<\tilde{b}<$ 1. Applying Coefficient Tightening on (10) sets the coefficient of $x_{r}$ to $\tilde{b}$. Subsequently, (10) propagates $x_{r}$ tightly.
@@
For pure binary constraints, a very competitive alternative to coefficient tightening in the reduction algorithm is based on Chvátal-Gomory cuts Elffers and Nordström (2018). In this reduction, relaxable variables with fractional coefficients in the reason constraint are weakened before applying Chvátal-Gomory rounding as in Definition 3. In Mexi et al. (2023) we show that applying the more general MIR cut instead yields a reduced reason constraint that is at least as strong.
Next, we present a further improved reduction technique also based on the MIR formula (4). The improvement comes from the fact that weakening becomes obsolete after complementing relaxable variables. We call this reduction cMIR-based reduction. In Section 5 we show one of the main results of this paper, which is the dominance of cMIR-based reduction over both Chvátal-Gomory and MIR-based reduction from Mexi et al. (2023). But first, we show that the reduced reason constraint from the cMIR-based reduction propagates $x_{r}$ tightly.
@@
Note that $10<\tilde{b}<$ 1 by Assumption 1. Indeed, since $C_{\textnormal{reason}}$ propagates non-tightly, we have that $\min\{b-\sum_{j\in J}a_{j}x_{j}\}$ is between 0 and 1. After applying MIR to the complemented constraint we obtain
@@
Complementing back gives us (11). Finally, we can use the fact that for $j\notin P$, $x_{j}$ is fixed to 0 and that $\psi(-a_{j})\leq 0$ to show that $C_{\textnormal{cMIR}}$ propagates $x_{r}$ tightly.. The propagated bound is
[…]
@@ 5 Dominance Relationships
The ultimate goal is to find a reduction technique that yields the strongest possible reduced constraint to use in the resolution step of conflict analysis. In this section, we discuss dominance relationships between the different reduction techniques in the following sense.
@@
The main goal of this section is to compare the cMIR-based reduction from Section 4.3 with the wMIR-based reduction from Mexi et al. (2023), which we recall next. Let $C_{\textnormal{reason}}:x_{r}+\sum_{j}a_{j}x_{j}\geq b$ be as in Assumption 1, propagating $x_{r}$ non-tightly. Then the reduced constraint
@@
Example 2.
Consider the following constraint
$C_{\textnormal{reason}}:\,\,3x_{1}+3x_{2}+3x_{3}+2x_{4}\geq 7$
which propagates $x_{4}$ to 1 non-tightly if $x_{1}$ is fixed to 00 at the local domain. By dividing $C_{\textnormal{reason}}$ by 2, weakening the relaxable variables $x_{2},x_{3}$ and applying MIR, we obtain the modified reduced constraint
$C_{\textnormal{wMIR}}:2x_{1}+x_{4}$ $\geq 1,$ (16)
which propagates $x_{4}$ tightly.
Our proposed cMIR-based reduction (11) can be used to obtain a stronger reason constraint. First, it complements $x_{2},x_{3}$ in $C_{\textnormal{reason}}$,
$3x_{1}-3\overline{x}_{2}-3\overline{x}_{3}+2x_{4}$ $\geq 1,$
then it divides $C_{\textnormal{reason}}$ by $d=2$ and applies MIR which yields the constraint
$2x_{1}-\overline{x}_{2}-\overline{x}_{3}+x_{4}$ $\geq 1,$
Complementing the variables $\overline{x}_{2},\overline{x}_{3}$ back to the original variable space yields
$C_{\textnormal{cMIR}}:2x_{1}+x_{2}+x_{3}+x_{4}$ $\geq 3,$ (17)
which propagates $x_{4}$ tightly. Obviously, (17) dominates (16) since weakening $x_{2},x_{3}$ in (17) gives exactly (16).
For completeness, we also recall that the coeffficient tightening-based reduction dominates the clausal-based reduction, and that the wMIR-based reduction dominates the Chvátal-Gomory-based reduction Mexi et al. (2023). Hence, Proposition 4 shows that the cMIR-based reduction also dominates the Chvátal-Gomory-based reduction. In Gocht et al. (2019) the authors show that the Chvátal-Gomory and coefficient tightening reduction algorithms are incomparable. Similar arguments can be used to show the same result for the MIR reduction from Proposition 3 and the coefficient tightening reduction from Algorithm 2. Details can be found in Mexi et al. (2023), Gocht et al. (2019).
@@ 6 Conflict Analysis for Mixed Binary Programs
Our next goal is to extend cut-based conflict analysis to the class of mixed binary programs. In Section 4 we have seen that for pure binary programs it is always possible to reduce the reason constraint to guarantee that in each iteration of conflict analysis the linear combination of the current conflict constraint and the reduced reason constraint remains infeasible under the local bounds, and hence the conflict analysis invariant is preserved. The key property of the reduced reason constraint is that it propagates the resolved variable tightly even when considered over the reals. In the case of MBPs, continuous variables are always propagated tightly. Hence, from Proposition 1 it follows that whenever resolving a continuous variable, the linear combination of the current conflict constraint and the reason constraint that propagated the variable we are resolving on remains infeasible under the local bounds.
@@
$C_{1}:$ $-2x_{1}-4y_{1}-2y_{2}$ $\geq-3$
$C_{2}:$ $20x_{1}+5y_{1}-y_{2}$ $\geq 4$
$C_{3}:$ $-20x_{1}+5y_{1}-10y_{2}$ $\geq-16$
$C_{4}:$ $-y_{2}-x_{2}$ $\geq 0$
$C_{5}:$ $y_{2}-x_{3}$ $\geq 0$
where $x_{1}$, $x_{2}$ and $x_{3}$ are binary variables, $y_{1}$ is a continuous variable with global bounds $[0,1]$, and $y_{2}$ is a continuous variable with global bounds $[-1,1]$. After branching on $x_{2}$, in the subproblem $x_{2}=0$ we can deduce $y_{2}\leq 0$ from $C_{4}$. Next, from $C_{5}$ we can deduce $x_{3}=0$ and $y_{2}\geq 0$. The constraint $C_{1}$ propagates $y_{1}\leq 3/4$ which leads to $C_{2}$ propagating $x_{1}\geq 1$. Under the local bounds, the constraint $C_{3}$ is infeasible. Eliminating $x_{1}$ with $C_{2}$ being the reason constraint and $C_{3}$ the conflict, we obtain $C_{2}+C_{3}:10y_{1}-11y_{2}\geq-12$ which is not infeasible under the local bounds.
The question that arises is whether we can construct a globally valid inequality that is violated in the local bounds by using only the constraint $C_{2}$ (reason) and $C_{3}$ (conflict) and the global bounds of the variables. We will show that not only it is impossible to do so by the reduction algorithm or the cutting planes we defined in Section 4, but also that no such inequality can be constructed if we do not consider more information about the problem.
To do so, we compute the mixed integer hull of the system consisting of the facets of the polyhedron defined by the constraints $C_{2}$ and $C_{3}$ and the global bounds of the variables. We compute the mixed integer hull by using the double description method Fukuda and Prodon (1995) implemented in Polymake Gawrilow and Joswig (2000). The facets are given by:
$F_{1}:$ $18x_{1}+25y_{1}-5y_{2}$ $\geq 20$
$F_{2}:$ $3x_{1}+5y_{1}$ $\geq 3$
$F_{3}:$ $-9x_{1}+5y_{1}-10y_{2}$ $\geq-5$
$F_{4}:$ $y_{1}$ $\leq 1$
$F_{5}:$ $x_{1}$ $\leq 1$
$F_{6}:$ $x_{1}$ $\geq 0$
$F_{7}:$ $y_{2}$ $\geq-1$
It is easy to verify that the facets $F_{1},\dots,F_{7}$ of the mixed integer hull are not violated under the local bounds. Moreover, the LP consisting of the facets and the local bounds is feasible. Therefore, we cannot find any globally valid inequality that is violated in the local bounds by only using the constraints $C_{2}$, $C_{3}$ and the global bounds of the variables.
@@ 6.1 Reduction Algorithm for Mixed Binary Programs
Comparing the right-hand sides of (18) and (20), we see that the linear combination (19) implies a bound at least as strong as $C_{\textnormal{reason}}$. ∎
@@
$00\geq(a_{r}+a^{\prime}_{r})x_{r}\geq b+b^{\prime}-\sum_{j\neq\{r,c\}}(a_{j}+a^{\prime}_{j})x_{j}\geq\kappa>$ 0 .
@@
Algorithm 3 shows the reduction algorithm for MBP. It takes as input the reason constraint $C_{\textnormal{reason}}$ propagating in state $\rho_{\hat{p},\hat{q}}$, the conflict constraint $C_{\textnormal{confl}}$ infeasible in state $\rho_{\bar{p},\bar{q}}$, and the binary variable to resolve $x_{r}$. Then it iteratively resolves all non-relaxable continuous variables from the reason constraint until no such variable exists. The set of non-relaxable continuous variables at a state $\rho_{p,q}$ is denoted by $\textnormal{nr}(C_{\textnormal{reason}},\rho_{p,q})\cap K$. The algorithm may terminate early if the reason constraint becomes infeasible under the local bounds (see Remark 3). Finally, after the main loop, we apply the reduction algorithm as in the binary case. The algorithm always terminates since the number of states is finite and it explores the states in a monotonically decreasing order with respect to the lexicographical order of the indices $(p,q)$ and the non-relaxable continuous variables appearing in the reason constraint.
@@
2$(\tilde{p},\tilde{q})\leftarrow(\hat{p},\hat{q})$
3 while $\textnormal{nr}(C_{\textnormal{reason}},\rho_{\tilde{p},\tilde{q}})\cap K\neq\varnothing$  do
4       $(\tilde{p},\tilde{q})\leftarrow\max\{(p,q)\,|\,\textrm{varidx}(\rho_{p,q})\in$ $\textnormal{nr}(C_{\textnormal{reason}},\rho_{p,q})\cap K,\,(p,q)\prec(\tilde{p},\tilde{q})\}$
5       $c\leftarrow\textrm{varidx}(\rho_{\tilde{p},\tilde{q}})$
6       $C_{\textnormal{cont}}\leftarrow\textrm{reason}(\rho_{\tilde{p},\tilde{q}})$

### v2

@@ 4 Reduction Techniques for Pure Binary Programs
Algorithm˜1 shows the base algorithm for all variants of cut-based conflict analysis considered in this paper. The algorithm is initialized with an infeasible state $\rho_{p,q}$ and a conflicting constraint $C_{\textnormal{confl}}$ in $\rho_{p,q}$. First, the learned conflict constraint $C_{\textnormal{learn}}$ is set to the conflict constraint $C_{\textnormal{confl}}$. In each iteration, the state $\rho_{p,q}$ is set to the smallest state, with respect to the lexicographic order, such that the learned conflict constraint is infeasible, and we extract the variable $x_{r}$ whose bound was changed in $\rho_{p,q}$. If the bound change was due to propagation of a constraint, then we extract the reason constraint $C_{\textnormal{reason}}$ that propagated $x_{r}$. In line 1 we “reduce” the reason constraint such that the resolvent, i.e., linear combination, of $C_{\textnormal{learn}}$ and the reduced reason $C_{\textnormal{reason}}$ (Line 1) that cancels $x_{r}$ remains infeasible. The learned conflict constraint is set to the resolvent, which can then be strengthened in Line 1 by, e.g., applying coefficient tightening as in SAT resolution conflict analysis. We continue this process until we reach an FUIP ($C_{\textnormal{learn}}$ is asserting) or $C_{\textnormal{learn}}$ proves global infeasibility.
@@ 10    $C_{\textnormal{learn}}\leftarrow\textnormal{strengthen}(C_{\textnormal{learn}})$
11
12
13return $C_{\textnormal{learn}}$
@@
The conflict analysis algorithm can fail if the resolvent becomes feasible. As explained in the previous section, this can occur only if the reason constraint $C_{\textnormal{reason}}$ does not propagate tightly. In the following, we present various reduction algorithms for the reason constraint $C_{\textnormal{reason}}$ that ensure that this constraint propagates tightly enough: the SAT-like clausal-based reduction Achterberg (2007a), which reduces the reason constraint to a clause responsible for the propagation; the coefficient tightening-based reduction, which is a generalization of Chai and Kuehlmann (2005); and a stronger version of the MIR-based reduction presented in Mexi et al. (2023).
@@
The reason constraint is of the form $C_{\textnormal{reason}}:x_{r}+\sum_{j\in J}a_{j}x_{j}\geq b,$ where $r\not\in J\subseteq\mathcal{I}$ and $a_{j}\geq 0$ for all $j\in J$. Furthermore, $C_{\textnormal{reason}}$ propagates $x_{r}$ in $\rho$ non-tightly, i.e., $b-\max\sum_{j\in J}a_{j}x_{j}\,\not\in\,\mathbb{Z}.$
@@ 4.1 Coefficient Tightening-based Reduction
algorithm˜2 summarizes the coefficient tightening-based reduction. Similar to the implementation in a PB solver Chai and Kuehlmann (2005), in each iteration, the algorithm picks a relaxable variable in the reason constraint different from the variable we are resolving on and weakens it. Then it applies coefficient tightening to the resulting constraint. After each iteration, the algorithm checks if the resolvent of the reason constraint and conflict constraint is infeasible. In this case, the algorithm terminates and returns the reduced reason constraint. A proof that the reduction algorithm produces a reason constraint that propagates strongly enough, in that the reduced reason is guaranteed to propagate tightly after all relaxable variables have been weakened, can be found in (Mexi et al. 2023).
Input : conflict constraint $C_{\textnormal{confl}}$, reason constraint $C_{\textnormal{reason}}$, variable to resolve $x_{r}$, state $\rho$
@@
[…]
@@ 5 $C_{\textnormal{reason}}\leftarrow\textnormal{coefTight}(C_{\textnormal{reason}})$
For pure binary constraints, a very competitive alternative to coefficient tightening in the reduction algorithm is based on Chvátal-Gomory cuts Elffers and Nordström (2018). In this reduction, relaxable variables with fractional coefficients in the reason constraint are weakened before applying Chvátal-Gomory rounding as in (2). In Mexi et al. (2023) we show that applying the more general MIR cut instead yields a reduced reason constraint that is at least as strong.
Next, we present a further improved reduction technique also based on the MIR formula (3). The improvement comes from the fact that weakening becomes obsolete after complementing relaxable variables. We call this reduction cMIR-based reduction. In Section˜5 we show one of the main results of this paper, which is the dominance of cMIR-based reduction over both Chvátal-Gomory and MIR-based reduction from Mexi et al. (2023). But first, we show that the reduced reason constraint from the cMIR-based reduction propagates $x_{r}$ tightly.
@@
Note that $0<\tilde{b}<1$ by Assumption 1Indeed, since $C_{\textnormal{reason}}$ propagates non-tightly, we have that $\min\{b-\sum_{j\in J}a_{j}x_{j}\}$ is between 0 and 1. After applying MIR to the complemented constraint we obtain
@@
Complementing back gives us (9). Finally, we can use the fact that for $j\notin P$, $x_{j}$ is fixed to 0 and that $\psi(-a_{j})\leq 0$ to show that $C_{\textnormal{cMIR}}$ propagates $x_{r}$ tightly. The propagated bound is
$1-\sum_{j\in P}\psi(-a_{j})-\max\Big\{$ $\sum_{j\notin P}\psi(a_{j})x_{j}-\sum_{j\in P}\psi(-a_{j})x_{j}\Big\}$
@@ 5 Dominance Relationships
A natural goal is to find a reduction technique that yields the strongest possible reduced constraint to use in the resolution step of conflict analysis. In this section, we discuss dominance relationships between the different reduction techniques in the following sense.
@@
The main goal of this section is to compare the cMIR-based reduction from Section˜4.2 with the wMIR-based reduction from Mexi et al. (2023), which we recall next. Let $C_{\textnormal{reason}}:x_{r}+\sum_{j}a_{j}x_{j}\geq b$ be as in Assumption 1, propagating $x_{r}$ non-tightly. Then the reduced constraint
@@
For completeness, we recall that the wMIR-based reduction dominates the Chvátal-Gomory-based reduction as described in Mexi et al. (2023). Hence, Proposition˜3 shows that the cMIR-based reduction also dominates the Chvátal-Gomory-based reduction. From Gocht et al. (2019) it follows that the Chvátal-Gomory and coefficient tightening reduction algorithms are incomparable. Similar arguments can be used to show the same result for the MIR reduction from Proposition˜2 and the coefficient tightening reduction from Algorithm˜2. Details can be found in Mexi et al. (2023), Gocht et al. (2019).
@@ 6 Conflict Analysis for Mixed Binary Programs
Our next goal is to extend cut-based conflict analysis to the class of mixed binary programs. In Section˜4 we have seen that for pure binary programs it is always possible to reduce the reason constraint to guarantee that in each iteration of conflict analysis the linear combination of the current conflict constraint and the reduced reason constraint remains infeasible under the local bounds, and hence the conflict analysis invariant is preserved. The key property of the reduced reason constraint is that it can be made to propagate the resolved variable tightly even when considered over the reals. In the case of MBPs, continuous variables are always propagated tightly. Hence, from Proposition˜1 it follows that whenever resolving a continuous variable, the linear combination of the current conflict constraint and the reason constraint that propagated the variable we are resolving on remains infeasible under the local bounds.
@@
$C_{1}:\,-2x_{1}-4y_{1}-2y_{2}\geq-3,\quad C_{2}:\,20x_{1}+5y_{1}-y_{2}\geq 4,\quad C_{3}:\,-20x_{1}+5y_{1}-10y_{2}\geq-16,\quad$
$C_{4}:\,-y_{2}-x_{2}\geq 0,\quad C_{5}:\,y_{2}-x_{3}\geq 0$
where $x_{1}$, $x_{2}$ and $x_{3}$ are binary variables, $y_{1}$ is a continuous variable with global bounds $[0,1]$, and $y_{2}$ is a continuous variable with global bounds [1,1][-1,1]. After branching on $x_{2}$, in the subproblem $x_{2}=0$ we can deduce $y_{2}\leq 0$ from $C_{4}$. Next, from $C_{5}$ we can deduce $x_{3}=0$ and $y_{2}\geq 0$. The constraint $C_{1}$ propagates $y_{1}\leq 3/4$ which leads to $C_{2}$ propagating $x_{1}\geq 1$. Under the local bounds, the constraint $C_{3}$ is infeasible. Eliminating $x_{1}$ with $C_{2}$ being the reason constraint and $C_{3}$ the conflict, we obtain $C_{2}+C_{3}:10y_{1}-11y_{2}\geq-12$ which is not infeasible under the local bounds. Next we show that it is not possible to find any globally valid inequality that is violated in the local bounds by considering only the constraints $C_{2}$ and $C_{3}$. Consider the two points $p_{1}=(1,0,0,0,-0.4)$ and $p_{2}=(0,0,0,0.88,0.4)$. Both points are integer feasible for the constraints $C_{2}$ and $C_{3}$. Since any globally valid inequality for the system $C_{2}$ and $C_{3}$ must satisfy the two points, it must also satisfy any convex combination $p^{\lambda}=\lambda p_{1}+(1-\lambda)p_{2}$ for $\lambda\in[0,1]$. However, the point $p^{0.5}=(0.5,0,0,0.44,0)$ lies in the local domain of the variables and hence cannot be separated by any globally valid inequality. Figure˜2 shows the non-empty intersection of the mixed integer hull with the local domain.
Figure 2: Left: Region defined by $C_{2}$, $C_{3}$, and the global domain; Middle: Mixed‐integer hull; Right: Mixed‐integer hull intersected with the local domain.
@@ 6.1 Reduction Algorithm for Mixed Binary Programs
Comparing the right-hand sides of (15) and (17), we see that the linear combination (16) implies a bound at least as strong as $C_{\textnormal{reason}}$.
@@
$0\geq(a_{r}+a^{\prime}_{r})x_{r}\geq b+b^{\prime}-\sum_{j\neq\{r,c\}}(a_{j}+a^{\prime}_{j})x_{j}\geq\kappa>0.$
@@
Algorithm˜3 shows the reduction algorithm for MBP. It takes as input the reason constraint $C_{\textnormal{reason}}$ propagating in state $\rho_{\hat{p},\hat{q}}$, the conflict constraint $C_{\textnormal{confl}}$ infeasible in state $\rho_{\bar{p},\bar{q}}$, and the binary variable to resolve $x_{r}$. Then it iteratively resolves all non-relaxable continuous variables from the reason constraint until no such variable exists. The set of non-relaxable continuous variables at a state $\rho_{p,q}$ is denoted by $\textnormal{nr}(C_{\textnormal{reason}},\rho_{p,q})$. The algorithm may terminate early if the reason constraint becomes infeasible under the local bounds (see Remark 3 in Appendix D). Finally, after the main loop, we apply the reduction algorithm as in the binary case. The algorithm always terminates since the number of states is finite and it explores the states in a monotonically decreasing order with respect to the lexicographical order of the indices $(p,q)$ and the non-relaxable continuous variables appearing in the reason constraint.
@@
[…]